The expression of the magnetic field associated with an electromagnetic wave in
vacuum is given by
B = (100T) y sin ( 2π × 10^8t + kz)
Determine the wave number, frequency and the direction of propagation of the wave,
and the magnitude and direction of the electric field associated with it.
1
Expert's answer
2020-03-09T10:58:36-0400

As per the given question,

"B = (100T) y sin ( 2\u03c0 \u00d7 10^8t + kz)"

Wave number is given = K

frequency = "10^8 Hz"

Direction of the propagation is in along the z axis

We know that the relation between the electric field (E)and the magnetic field(B) and the speed of light(c) is

"C=\dfrac{E}{B}"

"E=cB==3\times 10^{8}\times100y =3\times 10^{10}N/c"